OpenStudy (anonymous):

A bag contains 3 red marbles, 4 blue marbles, and 5 white marbles. What is the probability of pulling out a white marble, and then a blue marble if none are replaced?

Mehek (mehek14):

let's see we have 5 white marbles and a total of 12 marbles

Mehek (mehek14):

so the probability of getting a white marble is \(\dfrac{5}{12}\)

Mehek (mehek14):

when you take the white marble out and don't replace it, the total changes to 11

Mehek (mehek14):

you have 4 blue marbles and a total of 11 marbles \(\dfrac{4}{11}\)

Mehek (mehek14):

now to find out the probability of pulling out a white marble and then a blue marble is \(\dfrac{5}{12}*\dfrac{4}{11}\) @cadeneveah11 multiply that for me